# ResearchTown: Simulator of human research community

### v1

@@ 5 ResearchTown: Applying TextGNN to community graph
Stage 2: Paper writing. After paper reading, the next important research stage is paper writing. Different from paper reading, the paper writing process can be understood as inserting inserting a new data node to the community graph. Here, the new data node is non-existent before writing the paper, and the data node is created after the paper writing process. Concretely, by adapting Equation 4, the TextGNN layer for paper writing can be written as:
$\begin{split}\mathbf{h}_{v}&=\textsc{Agg}\big{(}\{f_{a}(\cdot),\mathbf{h}_{a}\mid(v,a)\in\mathcal{E}_{ad}\},\{\mathbf{h}_{d}\mid(v,d)\in\mathcal{E}_{dd}\}\big{)}\\ &=f_{g}\left(\left[\left\{f_{a}\big{(}\big{[}\mathbf{h}_{a},\mathbf{h}_{d}\big{]}\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\right\}\right]\right)\end{split}$ (6)
where $\mathbf{h}_{v}$ in Equation 4 is empty since the paper node content is non-existent before paper writing; here, $\mathcal{E}_{ad}$ specifically refers to authorship relations between agent and data nodes, and $\mathcal{E}_{dd}$ refers to citation relations within data nodes. A visualization of Equation 6 is shown in Figure 2 “Stage 2”.
Stage 3: Review writing. The review writing task is the final stage of the automatic research simulation, serving as a reflection stage in the multi-agent research simulator. The difference between the previous 2 stages is that first, the researchers involved during review writing are not the authors but the reviewers of the paper. Additionally, review writing is based on a written paper where $\mathbf{h}_{v}$ is no longer empty. Concretely, by adapting Equation 4, the TextGNN layer for review writing can be written as:
$\begin{split}\mathbf{r}_{v}&=\textsc{Agg}\big{(}\mathbf{h}_{v},\{f_{a}(\cdot),\mathbf{h}_{a}\mid(v,a)\in\mathcal{E}_{ad}\},\{\mathbf{h}_{d}\mid(v,d)\in\mathcal{E}_{dd}\}\big{)}\\ &=f_{g}\left(\left[\mathbf{h}_{v},\left\{f_{a}\big{(}\big{[}\mathbf{h}_{a},\mathbf{h}_{v},\mathbf{h}_{d}\big{]}\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\right\}\right]\right)\end{split}$ (7)
Summary: ResearchTown simulation algorithm. Utilizing the community graph $\mathcal{G}$, we propose a simulation algorithm named as ResearchTown. Overall, the simulation algorithm can be considered as a 2-layer GNN where the paper reading is the first layer of information aggregation. Both paper writing and review writing are the second layer of the GNN to generate the final simulation outputs. We formally summarize the research community simulation in Algorithm 1. To achieve better efficiency, the modified version for implementation is in Appendix §C.2.
@@
0:  Text-form paper $\mathbf{h}_{v}$ and review $\mathbf{r}_{v}$ for paper node $v$
@@
5:        $\mathbf{h}_{u}\leftarrow f_{u}\left(\left[\left\{\mathbf{x}_{d}\mid(u,d)\in\mathcal{E}_{ad}\right\}\right]\right)${Refer to Eq. (5)}
6:  $\mathbf{h}_{v}\leftarrow f_{g}\left(\left[\left\{f_{a}\big{(}\big{[}\mathbf{h}_{a},\mathbf{h}_{d}\big{]}\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\right\}\right]\right)$Refer to Eq. (6)}
7:  $\mathbf{r}_{v}\leftarrow f_{g}\left(\left[\mathbf{h}_{v},\left\{f_{a}\big{(}\big{[}\mathbf{h}_{a},\mathbf{h}_{v},\mathbf{h}_{d}\big{]}\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\right\}\right]\right)$Refer to Eq. (7)}
8:  return  $\mathbf{h}_{v}$, $\mathbf{r}_{v}$
@@ 6 Evaluating ResearchTown via masked node prediction task
Utilizing graph structures not only enables the development of the research simulation algorithm but also provides a natural way to evaluate it. As we show next, we propose to view research evaluation as a masked node prediction task, including evaluation for both paper writing and review writing.
@@
$\mathbf{h}_{v},\mathbf{r}_{v}=\textsc{ResearchTown}\left(\mathcal{G}(\mathcal{V},\mathcal{E});\{\mathbf{x}_{u}\mid u\in\mathcal{N}(v)\};v\right)$ (8)
where $\mathbf{h}_{v}$ is the text-form hidden states of a masked node $v$ and $\mathbf{r}_{v}$ is the text-form prediction output of a masked node $v$. Since we have real-world results for both paper writing and review, we treat them as ground truth even though they are not perfect because the goal of ResearchTown is to simulate the human research community rather than to find optimal solutions for papers and reviews ($\mathbf{h}_{v}^{*}$ for paper ground-truth and $\mathbf{r}_{v}^{*}$ for review ground-truth) and we can systematically evaluate both processes to check the effectiveness of our simulation algorithm. More specifically, since we can observe ground-truth papers $\mathbf{h}_{v}^{*}$ when evaluating the review quality, to avoid accumulated errors, we update Equation 7 during evaluation so that reviews $\mathbf{r}_{v}$ are generated based on $\mathbf{h}_{v}^{*}$, instead of $\mathbf{h}_{v}$:
$\mathbf{r}_{v}$ $=\textsc{Agg}\big{(}\mathbf{h}_{v}^{*},\{f_{a}(\cdot),\mathbf{h}_{a}\mid(v,a)\in\mathcal{E}_{ad}\},\{\mathbf{h}_{d}\mid(v,d)\in\mathcal{E}_{dd}\}\big{)}$ (9)
where $\textsc{sim}(\cdot,\cdot)$ represents an embedding-based similarity metric, such as voyage-3  [^2] https://blog.voyageai.com/2024/09/18/voyage-3/ and text-embedding-large-3  [^3] https://openai.com/index/new-embedding-models-and-api-updates/. By leveraging the LLM to generate structured embeddings for each question, this approach ensures a meaningful and consistent comparison of the generated and ground-truth papers. Details about the evaluation are available in Appendix §F.
where $\textsc{sim}(\cdot,\cdot)$ refers to the same similarity metric in paper writing evaluation. This metric emphasizes the recall rate of the generated review by measuring how well it captures the key points of the real-world review. Since each review consists of both strengths and weaknesses, we compute separate similarity scores for strengths and weaknesses. Additionally, since both $\mathbf{r}_{v}$ and $\mathbf{r}_{v}^{*}$ include a final score $\mathbf{s}_{v}$ and $\mathbf{s}^{*}_{v}$ as attributes, we calculate $\Delta\mathbf{s}_{v}=|\mathbf{s}_{v}-\mathbf{s}_{v}^{*}|$ to quantify the difference between the generated and real-world review scores. Details about the evaluation are available in Appendix §F.
@@ 7 Experimental settings
ResearchTown setting. We utilize gpt-4o-mini  [^4] We points to gpt-4o-mini-2024-07-18 for use. as the LLM backbone for implementing the agent functions, with the decoding temperature set to 00 to ensure reproducibility. To evaluate different aggregation strategies, we conduct experiments using specific types of edges connected to the target node: (1) Self-agg, where the aggregation relies solely on the target node; (2) Agent-agg, which includes the target node and its neighboring agent nodes; (3) Data-agg, which involves the target node and its neighboring data nodes; and (4) Global-agg, which incorporates the target node and all its neighboring nodes, including agent and data nodes. This experimental design enables a systematic comparison of the effects of different neighborhood information on the aggregation process. More details about different settings are available in Appendix §C.1.
@@
ResearchBench setting. To evaluate the effectiveness of ResearchTown for research simulation, we introduce ResearchBench, which consists of 1,000 paper writing tasks and 200 review writing tasks. All tasks are sourced from recent top-tier machine learning conferences such as NeurIPS 2024  [^5] https://neurips.cc/Conferences/2024 and ICLR 2024  [^6] https://openreview.net/group?id=ICLR.cc/2024/Conference. For paper writing tasks, we categorize them into three difficulty levels—hard (333 tasks), medium (334 tasks), and easy (333 tasks)—based on the similarity results of data-only aggregation. Specifically, for review writing tasks, the reviewers prepared for each paper are selected from the top 5 researchers most related to the paper, as reviewer information is not publicly available in the real world. More details about the data collection and prevention of information leakage during simulation are in Appendix §D.
8 Core results: In-distribution ResearchTown evaluation
In this section, we present the main results of our research simulation on ResearchBench. We evaluate existing paper nodes that have fully known neighborhoods, including their authors and citations, within the community graph. We refer to these scenarios as in-distribution cases.
Overall: ResearchTown can provide a realistic simulation of research activity. To evaluate research simulation, we utilize state-of-the-art embedding models including voyage-3 and text-embedding-3-large to compare the semantic similarity between simulated results and real-world results. For paper writing, as shown in Table 1, the overall similarity score obtained using text-embedding-3-large across 1,000 papers is 67.28. Notably, the score increases to 74.63 for an easy subset of the benchmark. These results demonstrate that paper writing with ResearchTown can produce realistic outputs compared to real-world ones. Moreover, it indicates that some ideas in top-tier conference papers are not hard to figure out and can be imagined by LLMs. For review writing, as shown in Table 2, the similarity scores are generally lower compared with paper writing, with strength-related scores averaging around 51 and weakness-related scores averaging around 47. This suggests that review writing is more challenging to generate with ResearchTown, particularly for weaknesses. A possible explanation is that real-world review data is often noisier and more diverse, especially for the weaknesses, making it harder to simulate accurately.
Special case: ResearchTown can potentially come up with high-impact ideas. Besides the full results on ResearchBench, we also evaluate ResearchTown under extreme conditions by attempting to simulate 100 of the most-cited machine learning papers from the past decade. As shown in Figure 4, ResearchTown achieves low similarity scores for papers introducing groundbreaking methods, such as “Layer Normalization” (Ba, 2016), or novel topics, such as “Energy and Policy Considerations for Deep Learning in NLP” (Strubell et al., 2019). However, the framework performs notably better on impactful papers focused on analysis or tool development. For instance, it achieves a similarity score exceeding 0.8 for papers like “Is BERT Really Robust? A Strong Baseline for Natural Language Attack on Text Classification and Entailment” (Jin et al., 2020), which provides adversarial analysis, and “Stanza: A Python Natural Language Processing Toolkit for Many Human Languages” (Qi et al., 2020), which offers a practical toolkit. These results suggest that generating high-impact research ideas may be more feasible than commonly perceived, and ResearchTown could potentially serve as a tool to inspire future impactful research.
@@ 8 Core results: In-distribution ResearchTown evaluation
Table 1: Embedding-based similarity scores for paper writing simulation. Similarity scores for paper writing are calculated based on Equation 10. “Hard”, “Medium”, and “Easy” correspond to three subsets of the paper writing tasks, while “Overall” refers to the performance across all parts. Detailed scores are shown in Appendix §G.
Experimental Setting text-embedding-large-3 $\uparrow$ voyage-3 $\uparrow$
Hard Medium Easy Overall Hard Medium Easy Overall
Self-agg 43.08 43.60 44.26 43.65 52.78 52.60 53.17 52.85
Agent-agg 52.32 54.77 56.75 54.61 57.05 58.77 60.39 58.74
Data-agg 55.83 67.39 76.93 66.72 60.57 69.69 78.14 69.47
Global-agg 59.59 67.50 74.74 67.28 63.34 69.78 76.19 69.77
Table 2: Embedding-based similarity score for review writing simulation. Similarity scores for both strengths and weaknesses of the reviews are calculated based on Equation 11. $\Delta\mathbf{s}$ refers to the average difference of review scores between ground-truth ones and generated ones.
Experimental Setting text-embedding-large-3 $\uparrow$ voyage-3 $\uparrow$ $\Delta\mathbf{s}$downarrow
Strength Weakness Strength Weakness
Self-agg 51.23 47.16 65.18 61.24 1.27
Agent-agg 51.66 46.75 66.03 61.29 1.19
Data-agg 51.45 47.62 65.57 61.74 1.26
Global-agg 51.51 47.17 66.01 61.39 1.55
Paper writing: multi-researchers participation improves paper quality. As shown in Table 1, cited papers contribute more effectively than authors in paper writing simulation, with data-aggregation achieving a score of 66.72 compared to 54.61 for agent-aggregation. The best results are obtained by combining both, surpassing data aggregation by 0.56 points. Researchers are particularly beneficial under hard scenarios, improving the text-embedding-large-3 score from 55.83 to 59.59, likely due to the inclusion of multi-hop paper information from researchers.
Review writing: multi-reviewers participation improves review quality. Unlike paper writing, review writing relies heavily on the target paper, making reviewers and cited papers less impactful, with differences limited to within 1 point. However, as shown in Table 2, adding additional information consistently improves performance over the self-aggregation baseline. Agent aggregation performs best for writing strengths and assigning scores, while data aggregation achieves the best results for writing weaknesses. This likely reflects the need for related work comparisons to identify weaknesses, while multiple reviewers contribute to a more comprehensive evaluation of strengths. Interestingly, combining agent and data aggregation leads to a decrease in score differences, possibly because the presence of related papers causes reviewers to apply stricter novelty judgments.
Figure 3: Similarity score distribution for paper writing simulation of 100 high-impact papers.
@@
Figure 4: Ablation study on the number of cited papers involved in paper writing simulation.
@@
Figure 5: Ablation study on the number of researchers involved in paper writing simulation.
@@
Figure 6: Ablation study on the number of reviewers involved in review writing simulation.

### v2

@@ 5 ResearchTown: Applying TextGNN to Community Graph
$\triangleright$ Stage 2: Paper writing. After paper reading, the next important research stage is paper writing. Different from paper reading, the paper writing process can be understood as inserting a new data node into the community graph. Here, the new data node is non-existent before writing the paper, and the data node is created after the paper writing process. Concretely, by adapting Equation 5, the TextGNN layer for paper writing can be written as:
$\mathbf{h}_{v}$ $=\textsc{AGG}\Big{(}\big{\{}f_{a}(\cdot),\big{\{}\mathbf{h}_{d}\mid(v,d)\in\mathcal{E}_{dd}\big{\}},\mathbf{h}_{a}\mid(v,a)\in\mathcal{E}_{ad}\big{\}}\Big{)}$ (7)
$=f_{g}\Big{(}\Big{[}\big{\{}f_{a}\big{(}[\mathbf{h}_{a},\mathbf{h}_{d}]\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\big{\}}\Big{]}\Big{)}$
where $\mathbf{h}_{v}$ in Equation 5 is empty since paper node contents are non-existent before paper writing; $\mathcal{E}_{ad}$ specifically refers to authorship relations between agent and data nodes, and $\mathcal{E}_{dd}$ refers to citation relations within data nodes. A visualization of Equation 7 is shown in Figure 2 “Stage 2”.
$\triangleright$ Stage 3: Review writing. The review writing task is the final stage of the automatic research simulation, serving as a reflection stage in the multi-agent research simulator. The difference between the previous 2 stages is that, first, the researchers involved during review writing are not the authors but the reviewers of the paper. Additionally, review writing is based on a written paper where $\mathbf{h}_{v}$ is no longer empty. Concretely, by adapting Equation 5, the TextGNN layer for review writing can be written as:
$\mathbf{r}_{v}$ $=\textsc{AGG}\Big{(}\mathbf{h}_{v},,\big{\{}\mathbf{h}_{d}\mid(v,d)\in\mathcal{E}_{dd}\big{\}}\big{\{}f_{a}(\cdot),\mathbf{h}_{a}\mid(v,a)\in\mathcal{E}_{ad}\big{\}}\Big{)}$ (8)
$=f_{g}\Big{(}\Big{[}\mathbf{h}_{v},\big{\{}f_{a}\big{(}[\mathbf{h}_{a},\mathbf{h}_{v},\mathbf{h}_{d}]\big{)}\mid(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\big{\}}\Big{]}\Big{)}$
$\triangleright$ Summary: ResearchTown simulation algorithm. Utilizing the community graph $\mathcal{G}$, we propose a simulation algorithm named as ResearchTown. Overall, the simulation algorithm can be considered as a 2-layer GNN where the paper reading is the first layer of information aggregation. Both paper writing and review writing are the second layer of the GNN to generate the final simulation outputs. We formally summarize the research community simulation in Algorithm 1. To achieve better efficiency, the modified version for implementation is in Appendix §D.2.
@@
0:  paper content $\mathbf{h}_{v}$ and review content $\mathbf{r}_{v}$ for paper node $v$
@@
5:       $\mathbf{h}_{u}\leftarrow f_{u}\left(\left[\left\{\mathbf{x}_{d}\mid(u,d)\in\mathcal{E}_{ad}\right\}\right]\right)$triangleright Eq. (6)
6:    end if
7:  end for
8:  $\mathbf{h}_{v}\leftarrow f_{g}\Big{(}\Big{[}\big{\{}f_{a}([\mathbf{h}_{a},\mathbf{h}_{d}])\mid$
9:    $(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\big{\}}\Big{]}\Big{)}$triangleright Eq. (7)
10:  $\mathbf{r}_{v}\leftarrow f_{g}\Big{(}\Big{[}\mathbf{h}_{v},\{f_{a}([\mathbf{h}_{a},\mathbf{h}_{v},\mathbf{h}_{d}])\mid$
11:    $(v,a)\in\mathcal{E}_{ad},(v,d)\in\mathcal{E}_{dd}\}\Big{]}\Big{)}$triangleright Eq. (8)
12:  return $\mathbf{h}_{v}$, $\mathbf{r}_{v}$
@@ 6 Evaluating ResearchTown via Masked Node Prediction Task
Utilizing graph structures not only enables the design of the research simulation algorithm but also provides a natural way to evaluate it. As we show next, we propose to view research evaluation as a masked node prediction task, including evaluation for both paper writing and review writing.
@@
$\begin{split}\mathbf{h}_{v},\mathbf{r}_{v}&=\textsc{ResearchTown}\Big{(}\mathcal{G}(\mathcal{V},\mathcal{E});\{\mathbf{x}_{u}\mid u\in\mathcal{N}(v)\};v\Big{)}\end{split}$ (9)
where $\mathbf{h}_{v}$ is the text-form hidden states of a masked node $v$ and $\mathbf{r}_{v}$ is the text-form prediction output of a masked node $v$. Since we have real-world results for both paper writing and review, we treat them as ground truth even though they are not perfect because the goal of ResearchTown is to simulate the human research community rather than to find optimal solutions for papers and reviews ($\mathbf{h}_{v}^{*}$ for paper ground-truth and $\mathbf{r}_{v}^{*}$ for review ground-truth) and we can systematically evaluate both processes to check the effectiveness of our simulation algorithm. More specifically, since we have access to ground-truth papers $\mathbf{h}_{v}^{*}$ when evaluating the review writing simulation, to avoid accumulated errors, we update Equation 8 during evaluation so that reviews $\mathbf{r}_{v}$ are generated based on $\mathbf{h}_{v}^{*}$, instead of $\mathbf{h}_{v}$:
$\begin{split}\mathbf{r}_{v}&=\textsc{AGG}\Big{(}\mathbf{h}_{v}^{*},\big{\{}\mathbf{h}_{d}\mid(v,d)\in\mathcal{E}_{dd}\big{\}},\big{\{}f_{a}(\cdot),\mathbf{h}_{a}\mid(v,a)\in\mathcal{E}_{ad}\big{\}}\Big{)}\end{split}$ (10)
Evaluation metric. We utilize state-of-the-art embedding models like text-embedding-large-3  [^1] https://openai.com/index/new-embedding-models-and-api-updates/ to build distance function for $d_{p}(\mathbf{h}_{v},\mathbf{h}_{v}^{*})$ and $d_{r}(\mathbf{r}_{v},\mathbf{r}_{v}^{*})$. More details related to formal embedding-based metric definitions for paper writing and review writing tasks are available in Appendix §F.
@@ 7 Experimental Settings
ResearchTown setting. We utilize GPT-4o-mini  [^2] We point to GPT-4o-mini-2024-07-18 for use. as the LLM backbone for implementing the agent functions, with the decoding temperature set to 00 to ensure reproducibility. To evaluate different aggregation strategies, we conduct experiments using specific types of nodes connected to the target node: (1) AGG-self, where the aggregation relies solely on the target node; (2) AGG-agent, which includes the target node and its neighboring agent nodes; (3) AGG-data, which involves the target node and its neighboring data nodes; and (4) AGG-global, which incorporates the target node and all its neighboring nodes, including agent and data nodes. We specifically refer to AGG-global as our proposed ResearchTown method for simulation, while the others serve as baselines. This experimental design enables a systematic comparison of the effects of different neighborhood information on the aggregation process. More details about different settings are available in Appendix §D.1.
@@
ResearchBench setting. To evaluate ResearchTown for research simulation, we introduce ResearchBench, which consists of 1,000 paper writing tasks and 200 review writing tasks. All tasks are sourced from recent top-tier machine learning conferences such as NeurIPS 2024  [^3] https://neurips.cc/Conferences/2024 and ICLR 2024  [^4] https://openreview.net/group?id=ICLR.cc/2024/Conference. Since most papers are released after the cutoff date of GPT-4o-mini, information leakage is not considered an issue. For paper writing tasks, we categorize them into three difficulty levels—hard (333 tasks), medium (334 tasks), and easy (333 tasks)—based on the similarity results of data-only aggregation. Specifically, for review writing tasks, the reviewers prepared for each paper are selected from the top 5 researchers most related to the paper, as reviewer information is not publicly available in the real world. More details about the data collection and prevention of information leakage during simulation are in Appendix §E.
@@
Table 1: Evaluation results for paper writing simulation. Hard, Medium, and Easy correspond to three subsets of the paper writing tasks with different difficulties, while Overall refers to the performance across all parts. Text-embedding-large-3 is used to build embedding-based similarity metrics. Comprehensive results are available in Appendix §J.
AGG Type Easy $\uparrow$ Medium $\uparrow$ Hard $\uparrow$ Overall $\uparrow$
AGG-self 46.42 45.92 45.90 46.08
AGG-agent 56.90 55.55 53.26 55.24
AGG-data 74.36 66.42 56.02 65.30
AGG-global 73.79 67.85 60.89 67.51
Table 2: Evaluation results for review writing simulation. For strength and weakness, it shows embedding-based similarity results. We use text-embedding-large-3 as embedding models and select 5 reviewers for running AGG-agent and AGG-global. $\Delta\mathbf{S}$ refers to the average difference of review scores between real-world ones and generated ones. $\bar{\mathbf{S}}$ refers to the average scores of generated ones. Comprehensive results are available in Appendix §J.
AGG Type Strength $\uparrow$ Weakness $\uparrow$ $\Delta\mathbf{S}$downarrow $\bar{\mathbf{S}}$
AGG-self 51.23 47.16 1.27 5.33
AGG-agent 51.66 46.75 1.19 5.40
AGG-data 51.45 47.62 1.26 5.30
AGG-global 51.51 47.17 1.55 5.00
@@
Figure 3: Ablation study on paper number. We select different sub-parts of cited papers in paper writing tasks.
@@
Figure 4: Ablation study on agent number. We select different numbers of agents for paper writing tasks.
@@
Figure 5: Ablation study on agent number. We select different numbers of agents for review writing tasks.
8 Core Results: In-distribution Evaluation
In this section, we present the main results of our research simulation on ResearchBench, including 1,000 paper writing tasks and 200 review writing tasks. We evaluate existing paper nodes that have fully known their content and their neighborhoods within the community graph. We refer to these scenarios as in-distribution cases.
Overall: ResearchTown can provide a realistic simulation of research activity. To evaluate research simulation, we utilize state-of-the-art embedding models (text-embedding-3-large) to compare the semantic similarity between simulated results and real-world results. For paper writing, as shown in Table 1, the overall similarity score obtained using text-embedding-3-large across 1,000 papers is 67.51. Notably, the score increases to 73.79 for an easy subset of the benchmark. These results demonstrate that paper writing with ResearchTown can produce realistic outputs compared to real-world ones. Moreover, it indicates that some ideas in top-tier conference papers are not hard to think of and can be imagined by LLMs. For review writing, as shown in Table 2, the similarity scores are generally lower compared with paper writing, with strength-related scores averaging around 51 and weakness-related scores averaging around 47. This suggests that review writing is more challenging to generate with ResearchTown, particularly for weakness identification. A possible explanation is that real-world review data is often noisier and more diverse, making it harder to simulate accurately.
Paper writing: participation of multi-researchers improves paper quality. As shown in Table 1, cited papers contribute more effectively than authors in the paper writing simulation, with data-aggregation achieving a score of 65.30 compared to 55.24 for agent-aggregation. The best results are obtained by combining both, surpassing data aggregation by 2.21 points. Researchers are particularly beneficial under difficult scenarios, improving the text-embedding-large-3 score from 56.02 to 60.89, likely due to the inclusion of multi-hop paper information from researchers.
Review writing: participation of multi-reviewers improves review quality. Unlike paper writing, review writing mainly relies on the paper that needs to be reviewed, making reviewers and cited papers less impactful, with differences limited to within 1 point. However, as shown in Table 2, adding additional information consistently improves performance over the self-aggregation baseline. Agent aggregation performs best for writing strengths and assigning scores, while data aggregation achieves the best results for writing weaknesses. This pattern likely reflects the role of related work comparisons in highlighting weaknesses, while multiple reviewers help provide a more balanced assessment of strengths. Interestingly, global aggregation leads to larger differences in scores. We consider it an exception since GPT-4o-mini tends to apply stricter novelty judgments under global aggregation—its average assigned score drops from 5.3 to 5.0. As shown in Table 3, this effect is not observed for Qwen-2.5-7B-Instruct or Deepseek-v3, which gain better results with global aggregation.